OpenStudy (anonymous):

y=-2x+5 change to standard form

OpenStudy (anonymous):

The standard form is \(ax+by+c=0\).

OpenStudy (anonymous):

I don't know how to get to my answer.

OpenStudy (anonymous):

Since the standard equation is \(=0\), then you must add 2x+5 on both sides.

OpenStudy (anonymous):

Can you do it?

OpenStudy (koikkara):

Given -2x+y=5 \(OR\) subtract -2x from each side @yoda5657 you may get the solution in \(y-y1=m(x-x1)\) form.